Bet'R Bilt Bikes just announced that their annual dividend for this coming year will be $1.6 a share and that all future dividends are expected to increase by 2% annually. What is the market rate of return if this stock is currently selling for $20 a share?
Answer #1

Stock price = Expected dividend/ (Rate of return - growth)

20 = 1.6/ (R-2%)

R-2% = 1.6/20

R = 0.08+ 0.02

Market rate of return = 10%